The vote count for candidate A was 7,310. The vote count for candidate B was 14,190. What percent of the voters voted for candidate A?
Found 2 solutions by MRperkins, gonzo:
Answer by MRperkins(300)

find the total number of voters. Then take (vote count for candidate A)/(total votes). This will give you the fraction and multiply this by 100 to get your percentage.
.
I hope this helps!
.
Private tutoring is available. Click on my name to go to my website or email me at justin.sheppard.tech@hotmail.com for more information. If you have any other questions you can direct them to me personally and I will answer them for you.

Answer by gonzo(654)

total vote count is 7310 + 14190 = 21500
percent of voters who voted for candidate A is:
7310 / 21500 = .34 * 100% = 34%
-----